There are two questions:
Can you tell me step by step what is 9(5x-2)= 45
The second problem is 10(3x+2)=70

Can you please step by step and the answer

Answer by longjonsilver(2297)

1. 9(5x-2)= 45

5x-2 = 45/9
5x-2 = 5
5x = 7
x = 7/5

2. 10(3x+2)=70

3x+2 = 70/10
3x+2 = 7
3x = 5
x = 5/3

jon.